OpenStudy (anonymous):

x ^{2} - 1 =8 is a _________ equation?

OpenStudy (anonymous):

Quadratic?

myininaya (myininaya):

yep

OpenStudy (anonymous):

How do you actually know the degree of an equation?

myininaya (myininaya):

the highest exponent

OpenStudy (anonymous):

of a polynomial?

OpenStudy (anonymous):

of a polynomial in the equation?

myininaya (myininaya):

yes polynomial=0 the degree of the polynomial is the highest exponent that can be found in that polynomial

OpenStudy (anonymous):

Ok , so when x^3 -1 = 8 it is a cubic ? But what if, x^3 + x + constant = 0? it is still cubic?

myininaya (myininaya):

yep yep

OpenStudy (anonymous):

so the highest exponent is considered. no matter how many other variables are there in the equation?

myininaya (myininaya):

yep yep

myininaya (myininaya):

of course we are talking about polynomials=0 right?

myininaya (myininaya):

I don't think other type of equations have special names

myininaya (myininaya):

x-1=4 this is a linear equation

myininaya (myininaya):

x-1-4=0 x-5=0

OpenStudy (anonymous):

thanks inaya.

myininaya (myininaya):

np